OpenStudy (anonymous):

Two moles of an ideal gas at ATM and 10degrees C is heated up to 150 degrees C. If the volume is held constant during this heating, what is the final pressure?

OpenStudy (anonymous):

you can use \[PV=nRT\] for this one too - since the two volumes are equal to each other, you can rearrange with equation to say \[V_1=\frac{n_1RT_1}{P_1}\] \[V_2=\frac{n_2RT_2}{P_2}\] \[V_1=V_2\] \[\frac{n_1RT_1}{P_1}=\frac{n_2RT_2}{P_2}\] Then solve for P2

OpenStudy (anonymous):

Thank you so much